# A Survey on Large Language Models for Code Generation

### v1

@@ 4. Large Langauge Models for Code Generation
Large language models (LLMs) with Transformer architecture have revolutionized a multitude of fields, and their application in code generation has been particularly impactful. These models follow a comprehensive process that starts with the curation and synthesis of code data, followed by a structured training approach that includes pre-training and fine-tuning, and the use of sophisticated prompt engineering techniques. Recent advancements have seen the integration of repository-level and retrieval-augmented code generation, as well as the development of autonomous coding agents. Furthermore, the evaluation of coding abilities of LLMs has become a critical component of this research area.
In the forthcoming sections, we will explore these dimensions of LLMs in the context of code generation in detail. Section 4.1 will address the data curation and processing strategies employed throughout the various stages of LLM development. Section 4.2 will discuss data synthesis methods designed to mitigate the scarcity of high-quality data. Section 4.3 will outline the prevalent model architectures used in LLMs for code generation. Moving to Section 4.4, we will examine the techniques for full parameter fine-tuning and parameter-efficient fine-tuning, which are essential for tailoring LLMs to code generation task. Section 4.5 will shed light on enhancing code quality through reinforcement learning, utilizing the power of feedback. Section 4.6 will delve into the strategic use of prompts to maximize the coding capabilities of LLMs. The innovative approaches of repository-level and retrieval-augmented code generation will be elaborated in Sections 4.7 and 4.8, respectively. Additionally, Section 4.9 will discuss the exciting field of autonomous coding agents. Lastly, Section 4.11 will provide insights into some of the practical applications that leverage LLMs for code generation, demonstrating the real-world impact of these sophisticated models. Through this comprehensive exploration, we aim to highlight the significance and potential of LLMs within the domain of automated code generation.
@@ 4.1. Data Curation & Processing
The exceptional performance of Large Language Models (LLMs) can be attributed to their training on large-scale and diverse datasets (Zan et al., 2023). Meanwhile, the extensive parameters of these models necessitate substantial data to unlock their full potential, in alignment with established scaling law (Kaplan et al., 2020; Hoffmann et al., 2022). For a general-purpose LLM, amassing a large-scale corpus of natural language from a variety of sources is imperative. Such sources include webpages, conversation data, books and news, scientific data, and code (Brown et al., 2020; Chowdhery et al., 2023; Bai et al., 2023; Touvron et al., 2023a, b; Yoo et al., 2024), while these data are often crawled from the web and must undergo meticulous and aggressive pre-processing (Raffel et al., 2020; Zhang et al., 2023b). Fortunately, multiple platforms and websites offer large-scale, open-source, and permissively licensed code corpora, such as GitHub [^7] https://github.com and Stack Overflow [^8] https://stackoverflow.com. Notably, the number of stars or forks of GitHub repositories has emerged as a valuable metric for filtering high-quality code datasets. In a similar vein, the quantity of votes on Stack Overflow can serve to discern the most relevant and superior answers.
@@
Figure 4. A diagram depicting the standard data preprocessing workflow utilized in the pre-training phase of large language models (LLMs) for code generation.
@@ 4.1.1. Pre-training
The remarkable success of bidirectional pre-trained language models (PLMs) such as BERT (Devlin et al., 2018) and unidirectional PLMs like GPT (Radford et al., 2018) has firmly established the practice of pre-training on large-scale unlabeled datasets to endow models with a broad spectrum of general knowledge. Extending this principle to the realm of code generation enables Large Language Models (LLMs) to assimilate fundamental coding principles, including the understanding of code structure dependencies, the semantics of code identifiers, and the intrinsic logic of code sequences (Chen et al., 2021; Wang et al., 2021; Guo et al., 2022; Wang et al., 2023b). In light of this advancement, there has been a proliferation of large-scale unlabeled code datasets proposed to serve as the foundational training ground for LLMs to develop coding proficiency. A brief introduction of these datasets is as follows, with the statistics available in Table 1.
@@
Table 1. The statistics of some commonly-used pre-training datasets for large language models (LLMs) aimed at code generation. The column labeled ‘#PL’ indicates the number of programming languages included in each dataset. It should be noted that in the CodeSearchNet (Husain et al., 2019) dataset, each file represents a function, and for the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022) datasets, only the code components are considered.
@@ 4.1.2. Instruction Tuning
Instruction tuning refers to the process of fine-tuning large language models (LLMs) using a collection of datasets that are structured as instructions. This method has demonstrated a considerable improvement in model performance and an enhanced ability to generalize to unseen tasks that the model has not previously encountered, as evidenced by recent studies (Ouyang et al., 2022; Chung et al., 2024). Leveraging the benefits of instruction tuning, instruction tuning has been expanded into coding domains, especially for code generation, which involves the automatic generation of the intended code from a natural language description. The promise of instruction tuning in this area has led numerous researchers to develop large-scale instruction-tuning datasets tailored for code generation. Below, we provide an overview of several notable datasets tailored for instruction tuning, with their respective statistics detailed in Table 2.
@@
  • Magicoder-OSS-Instruct-75k (Wei et al., 2023): is a 75k synthetic data generated through OSS-Instruct with gpt-3.5-turbo-1106 and used to train both Magicoder and Magicoder-S series models.
  • Self-OSS-Instruct-SC2-Exec-Filter-50k (Yuxiang Wei, 2024): Self-OSS-Instruct-SC2-Exec-Filter-50k is generated by StarCoder2-15B using the OSS-Instruct (Wei et al., 2023) data synthesis approach. It was subsequently used to fine-tune StarCoder-15B without any human annotations or distilled data from huge and proprietary LLMs.
Table 2. The statistics of several representative datasets used in instruction-tuning large language models (LLMs) for code generation. The column labeled ‘#PL’ indicates the number of programming languages encompassed by each dataset.
@@
Magicoder-OSS-Instruct-75k (Wei et al., 2023) 75k
@@ 4.1.3. Benchmarks
To rigorously assess the efficacy of Large Language Models (LLMs) for code generation, the research community has introduced a variety of high-quality benchmarks in recent years. Building on the foundational work by (Chen et al., 2021), numerous variations of the HumanEval dataset and additional benchmarks have emerged, aiming to evaluate a broader spectrum of code generation capabilities in LLMs. We roughly divide these benchmarks into six distinct categories based on their application contexts, including general-purpose, competitive programming, data science, multilingual, logical reasoning, and repository-level. The statistics for these benchmarks are presented in Table 3.
@@
  • HumanEval+ (Liu et al., 2024b): HumanEval+ extends the original HumanEval (Chen et al., 2021) benchmark by increasing the scale of the test cases by 80 times. As the test cases increase, HumanEval+ can catch significant amounts of previously undetected incorrect code synthesized by LLMs.
@@
  • MBPP+ (Liu et al., 2024b): MBPP+ enhances MBPP (Austin et al., 2021) by eliminating ill-formed problems and rectifying problems with incorrect implementations. The test scale of MBPP+ is also expanded by 35 times for test augmentation.
@@
  • ODEX (Wang et al., 2022d): ODEX is an open-domain dataset focused on the execution-based generation of Python code from natural language. It features 945 pairs of natural language queries and their corresponding Python code, all extracted from StackOverflow forums.
@@
  • ReCode (Wang et al., 2022c): Recode serves as a comprehensive robustness evaluation benchmark. ReCode applies perturbations to docstrings, function and variable names, code syntax, and code format, thereby providing multifaceted assessments of a model’s robustness performance.
@@
  • CodeContests (Li et al., 2022a): is a competitive programming dataset consisting of samples from various sources including Aizu, AtCoder, CodeChef, Codeforces, and HackerEarth. The dataset encompasses programming problems accompanied by test cases in the form of paired inputs and outputs, along with both correct and incorrect human solutions in multiple programming languages.
@@
Table 3. The detailed statistics of commonly-used benchmarks used in evaluating large language models (LLMs) for code generation. The column labeled ‘#PL’ indicates the number of programming languages included in each dataset. For the sake of brevity, we list the programming languages (PLs) for benchmarks that support fewer than or include five PLs. For benchmarks with six or more PLs, we provide only a numerical count of the PLs supported.
@@
HumanEval+ (Liu et al., 2024b) 164 Python 2023-05 https://huggingface.co/datasets/evalplus/humanevalplus
@@
MBPP+ (Liu et al., 2024b) 378 Python 2023-05 https://huggingface.co/datasets/evalplus/mbppplus
@@
ODEX (Wang et al., 2022d) 945 Python 2022-12 https://huggingface.co/datasets/neulab/odex
@@
ReCode (Wang et al., 2022c) 1,138 Python 2022-12 https://github.com/amazon-science/recode
@@
Repobench (Liu et al., 2023b) 27k Python, Java 2023-01 https://github.com/Leolty/repobench
@@
  • Repobench (Liu et al., 2023b): Repobench is a benchmark specifically used for evaluating repository-level code auto-completion systems. Supporting both Python and Java, it consists of three interconnected evaluation tasks: RepoBench-R (Retrieval), RepoBench-C (Code Completion), and RepoBench-P (Pipeline).
@@ 4.2. Data Synthesis
Numerous studies have demonstrated that high-quality datasets are integral to enhancing the performance of large language models (LLMs) in various downstream tasks (Brown et al., 2020; Meng et al., 2022; Xie et al., 2023; Zhou et al., 2024; Köpf et al., 2024; Wettig et al., 2024). For instance, the LIMA model, a 65B parameter LLaMa language model fine-tuned with a standard supervised loss on a mere 1,000 meticulously curated prompts and responses, achieved performance on par with, or even superior to, GPT-4 in 43% of evaluated cases. This figure rose to 58% when compared to Bard and 65% against DaVinci003, all without the use of reinforcement learning or human preference modeling (Zhou et al., 2024). The QuRating initiative strategically selects pre-training data embodying four key textual qualities — writing style, facts & trivia, required expertise, and educational value — that resonate with human intuition. Training a 1.3B parameter model on such data resulted in reduced perplexity and stronger in-context learning compared to baseline models (Wettig et al., 2024).
Despite these advancements, acquiring quality data remains a significant challenge due to issues such as data scarcity, privacy concerns, and prohibitive costs (Wang et al., 2023a; Liu et al., 2024a). Human-generated data is often labor-intensive and expensive to produce, and it may lack the necessary scope and detail to navigate complex, rare, or ambiguous scenarios. As a resolution to these challenges, synthetic data has emerged as a viable alternative. By generating artificial datasets that replicate the intricacies of real-world information, models such as GPT-3.5-turbo (OpenAI, 2022) and GPT-4 (Achiam et al., 2023) have enabled the creation of rich datasets without the need for human annotation (Wang et al., 2023a; Hämäläinen et al., 2023; Liu et al., 2024a; Laurer, 2024). This approach is particularly beneficial in enhancing the instruction-following capabilities of LLMs, with a focus on generating synthetic instruction-based data.
@@
Drawing on the successes of data synthesis for general-purpose Large Language Models (LLMs), researchers have expanded the application of synthetic data to the realm of code generation. The Code Alpaca model, as described in (Chaudhary, 2023), has been fine-tuned on a 7B and 13B LLaMA model using a dataset of 20k instruction-following examples for code generation. This dataset was created by text-davinci-003101010https://platform.openai.com and employed the Self-Instruct technique (Wang et al., 2023a). Building on this, the WizardCoder 15B (Luo et al., 2023) utilizes the Evol-Instruct technique to create an enhanced dataset of 78k evolved code instruction examples. This dataset originates from the initial 20k instruction-following dataset used by Code Alpaca (Chaudhary, 2023), which was also generated by text-davinci-003. The WizardCoder model, fine-tuned on the StarCoder (Li et al., 2023a) base model, achieved a 57.3% pass@1 on the HumanEval benchmarks. This performance not only surpasses all other open-source Code LLMs by a significant margin but also outperforms leading closed LLMs such as Anthropic’s Claude and Google’s Bard. In a similar vein, Magicoder (Wei et al., 2023) introduces a novel data synthesis approach termed OSS-INSTRUCT which enlightens LLMs with open-source code snippets to generate high-quality instruction data for coding tasks. It aims to address the inherent biases often present in synthetic data produced by LLMs. Building upon CodeLlama (Roziere et al., 2023), the MagicoderS-CL-7B model — fine-tuned with 75k synthetic instruction data using the OSS-INSTRUCT technique and with gpt-3.5-turbo-1106 as the data generator — has outperformed the prominent ChatGPT on the HumanEval Plus benchmark, achieving pass@1 of 66.5% versus 65.9%. In a noteworthy development, Microsoft has introduced the phi-1 model (Gunasekar et al., 2023), a more compact LLM of only 1.3B parameters. Despite its smaller size, phi-1 has been trained on high-quality textbook data sourced from the web (comprising 6 billion tokens) and supplemented with synthetic textbooks and exercises generated with GPT-3.5 (1 billion tokens). It has achieved pass@1 of 50.6% on HumanEval and 55.5% on MBPP, setting a new state-of-the-art for Python coding performance among existing small language models (SLMs). The latest contribution to this field is from the BigCode team, which has presented StarCoder2-15B-instruct (Yuxiang Wei, 2024), the first entirely self-aligned code LLM trained with a transparent and permissive pipeline. This model aligns closely with the OSS-INSTRUCT principles established by Magicoder, generating instructions based on seed functions filtered from the Stack v1 dataset (Kocetkov et al., 2022) and producing responses through self-validation. Unlike Magicoder, StarCoder2-15B-instruct employs its base model, StarCoder2-15B, as the data generator, thus avoiding reliance on large and proprietary LLMs like GPT-3.5-turbo (OpenAI, 2022).

### v2

@@ 5. Large Langauge Models for Code Generation
LLMs with Transformer architecture have revolutionized a multitude of fields, and their application in code generation has been particularly impactful. These models follow a comprehensive process that starts with the curation and synthesis of code data, followed by a structured training approach that includes pre-training and fine-tuning (instruction tuning), reinforcement learning with various feedback, and the use of sophisticated prompt engineering techniques. Recent advancements have seen the integration of repository-level and retrieval-augmented code generation, as well as the development of autonomous coding agents. Furthermore, the evaluation of coding abilities of LLMs has become a critical component of this research area. Figure 5 illustrates the general training, inference, and evaluation workflow for Code LLMs and their associated databases.
In the forthcoming sections, we will explore these dimensions of LLMs in the context of code generation in detail. Section 5.1 will address the data curation and processing strategies employed throughout the various stages of LLM development. Section 5.2 will discuss data synthesis methods designed to mitigate the scarcity of high-quality data. Section 5.3 will outline the prevalent model architectures used in LLMs for code generation. Moving to Section 5.4, we will examine the techniques for full parameter fine-tuning and parameter-efficient fine-tuning, which are essential for tailoring LLMs to code generation task. Section 5.5 will shed light on enhancing code quality through reinforcement learning, utilizing the power of feedback. Section 5.6 will delve into the strategic use of prompts to maximize the coding capabilities of LLMs. The innovative approaches of repository-level and retrieval-augmented code generation will be elaborated in Sections 5.7 and 5.8, respectively. Additionally, Section 5.9 will discuss the exciting field of autonomous coding agents. Section 5.10 discusses various evaluation strategies and offer an empirical comparison using the widely recognized HumanEval, MBPP, and the more practical and challenging BigCodeBench benchmarks to highlight the progressive enhancements in LLM capabilities for code generation. Furthermore, the ethical implications and the environmental impact of using LLMs for code generation are discussed in Section 5.11, aiming to establish a trustworthiness, responsibility, safety, efficiency, and green of LLM for code generation. Lastly, Section 5.12 will provide insights into some of the practical applications that leverage LLMs for code generation, demonstrating the real-world impact of these sophisticated models. Through this comprehensive exploration, we aim to highlight the significance and potential of LLMs within the domain of automated code generation.
@@ 5.1. Data Curation & Processing
The exceptional performance of LLMs can be attributed to their training on large-scale and diverse datasets (Zan et al., 2023). Meanwhile, the extensive parameters of these models necessitate substantial data to unlock their full potential, in alignment with established scaling law (Kaplan et al., 2020; Hoffmann et al., 2022). For a general-purpose LLM, amassing a large-scale corpus of natural language from a variety of sources is imperative. Such sources include webpages, conversation data, books and news, scientific data, and code (Brown, 2020; Chowdhery et al., 2023; Bai et al., 2023; Touvron et al., 2023a, b; Yoo et al., 2024), while these data are often crawled from the web and must undergo meticulous and aggressive pre-processing (Raffel et al., 2020; Zhang et al., 2023b). Fortunately, multiple platforms and websites offer large-scale, open-source, and permissively licensed code corpora, such as GitHub [^7] https://github.com and Stack Overflow [^8] https://stackoverflow.com. Notably, the number of stars or forks of GitHub repositories has emerged as a valuable metric for filtering high-quality code datasets. In a similar vein, the quantity of votes on Stack Overflow can serve to discern the most relevant and superior answers.
@@
Figure 7. A diagram depicting the standard data preprocessing workflow utilized in the pre-training phase of LLMs for code generation.
@@ 5.1.1. Pre-training
The remarkable success of bidirectional pre-trained language models (PLMs) such as BERT (Devlin et al., 2018) and unidirectional PLMs like GPT (Radford et al., 2018) has firmly established the practice of pre-training on large-scale unlabeled datasets to endow models with a broad spectrum of general knowledge. Extending this principle to the realm of code generation enables LLMs to assimilate fundamental coding principles, including the understanding of code structure dependencies, the semantics of code identifiers, and the intrinsic logic of code sequences (Chen et al., 2021; Wang et al., 2021; Guo et al., 2022; Wang et al., 2023b). In light of this advancement, there has been a proliferation of large-scale unlabeled code datasets proposed to serve as the foundational training ground for LLMs to develop coding proficiency. A brief introduction of these datasets is as follows, with the statistics available in Table 4.
@@
Table 4. The statistics of some commonly-used pre-training datasets for LLMs aimed at code generation. The column labeled ‘#PL’ indicates the number of programming languages included in each dataset. It should be noted that in the CodeSearchNet (Husain et al., 2019) dataset, each file represents a function, and for the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022) datasets, only the code components are considered.
@@ 5.1.2. Instruction Tuning
Instruction tuning refers to the process of supervised fine-tuning LLMs using a collection of datasets structured as various instructions, with the purpose of following a wide range of task instructions (Wei et al., 2021; Sanh et al., 2022; Ouyang et al., 2022; Chung et al., 2024). This method has demonstrated a considerable improvement in model performance and an enhanced ability to generalize to unseen tasks that the model has not previously encountered, as evidenced by recent studies (Ouyang et al., 2022; Chung et al., 2024). Leveraging the benefits of instruction tuning, instruction tuning has been expanded into coding domains, especially for code generation, which involves the automatic generation of the intended code from a natural language description. The promise of instruction tuning in this area has led numerous researchers to develop large-scale instruction-tuning datasets tailored for code generation. Below, we provide an overview of several notable datasets tailored for instruction tuning, with their respective statistics detailed in Table 5.
@@
  • Magicoder-OSS-Instruct-75k (Wei et al., 2023b): is a 75k synthetic data generated through OSS-Instruct with gpt-3.5-turbo-1106 and used to train both Magicoder and Magicoder-S series models.
  • Self-OSS-Instruct-SC2-Exec-Filter-50k (Yuxiang Wei, 2024): Self-OSS-Instruct-SC2-Exec-Filter-50k is generated by StarCoder2-15B using the OSS-Instruct (Wei et al., 2023b) data synthesis approach. It was subsequently used to fine-tune StarCoder-15B without any human annotations or distilled data from huge and proprietary LLMs.
Table 5. The statistics of several representative datasets used in instruction-tuning LLMs for code generation. The column labeled ‘#PL’ indicates the number of programming languages encompassed by each dataset.
@@
evol-codealpaca-v1 (theblackcat102, 2023) 110K - 2023-07 https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1
Magicoder-OSS-Instruct-75k (Wei et al., 2023b) 75k
@@ 5.1.3. Benchmarks
To rigorously assess the efficacy of LLMs for code generation, the research community has introduced a variety of high-quality benchmarks in recent years. Building on the foundational work by (Chen et al., 2021), numerous variations of the HumanEval dataset and additional benchmarks have emerged, aiming to evaluate a broader spectrum of code generation capabilities in LLMs. We roughly divide these benchmarks into six distinct categories based on their application contexts, including general-purpose, competitive programming, data science, multilingual, logical reasoning, and repository-level. It is important to highlight that logical reasoning encompasses math-related benchmarks, as it aims to create “code-based solutions” for solving complex mathematical problems (Chen et al., 2022b; Gao et al., 2023a; Zhou et al., 2023a). This strategy can therefore mitigate the limitations of LLMs in performing intricate mathematical computations. The statistics for these benchmarks are presented in Table 6.
@@
  • HumanEval+ (Liu et al., 2024c): HumanEval+ extends the original HumanEval (Chen et al., 2021) benchmark by increasing the scale of the test cases by 80 times. As the test cases increase, HumanEval+ can catch significant amounts of previously undetected incorrect code synthesized by LLMs.
@@
  • MBPP+ (Liu et al., 2024c): MBPP+ enhances MBPP (Austin et al., 2021) by eliminating ill-formed problems and rectifying problems with incorrect implementations. The test scale of MBPP+ is also expanded by 35 times for test augmentation.
@@
  • ODEX (Wang et al., 2022e): ODEX is an open-domain dataset focused on the execution-based generation of Python code from natural language. It features 945 pairs of natural language queries and their corresponding Python code, all extracted from StackOverflow forums.
@@
  • ReCode (Wang et al., 2022d): Recode serves as a comprehensive robustness evaluation benchmark. ReCode applies perturbations to docstrings, function and variable names, code syntax, and code format, thereby providing multifaceted assessments of a model’s robustness performance.
@@
  • BigCodeBench (Zhuo et al., 2024a): BigCodeBench has 1,140 complex Python programming tasks, covering 723 function calls from 139 popular libraries across 7 domains. This benchmark is specifically designed to assess LLMs’ ability to call multiple functions from cross-domain libraries and follow complex instructions to solve programming tasks, helping to bridge the evaluation gap between isolated coding exercises and the real-world programming scenario.
  • ClassEval (Du et al., 2024): ClassEval is a manually-crafted benchmark consisting of 100 classes and 412 methods for evaluating LLMs in the class-level code generation scenario. Particularly, the task samples of ClassEval present higher complexities, involving long code generation and sophisticated docstring information, thereby benefiting the evaluation of the LLMs’ capabilities in generating complicated code.
  • NaturalCodeBench (Zhang et al., 2024c): NaturalCodeBench is a comprehensive code benchmark featuring 402 high-quality problems in Python and Java. These problems are selected from natural user queries from online coding services and span 6 distinct domains, shaping an evaluation environment aligned with real-world applications.
@@
  • CodeContests (Li et al., 2022a): CodeContests is a competitive programming dataset consisting of samples from various sources including Aizu, AtCoder, CodeChef, Codeforces, and HackerEarth. The dataset encompasses programming problems accompanied by test cases in the form of paired inputs and outputs, along with both correct and incorrect human solutions in multiple programming languages.
  • LiveCodeBench (Naman Jain et al., 2024): LiveCodeBench is a comprehensive and contamination-free benchmark for evaluating a wide array of code-related capabilities of LLMs, including code generation, self-repair, code execution, and test output prediction. It continuously gathers new coding problems from contests across three reputable competition platforms: LeetCode, AtCoder, and CodeForces. The latest release of the dataset includes 713 problems that were released between May 2023 and September 2024.
@@
  • CRUXEval (Gu et al., 2024): CRUXEval contains 800 Python functions, each paired with an input-output example. This benchmark supports two tasks: input prediction and output prediction, designed to evaluate the code reasoning, understanding, and execution capabilities of code LLMs.
Table 6. The detailed statistics of commonly-used benchmarks used in evaluating LLMs for code generation. The column labeled ‘#PL’ indicates the number of programming languages included in each dataset. For the sake of brevity, we list the programming languages (PLs) for benchmarks that support fewer than or include five PLs. For benchmarks with six or more PLs, we provide only a numerical count of the PLs supported.
@@
HumanEval+ (Liu et al., 2024c) 164 Python 2023-05 https://huggingface.co/datasets/evalplus/humanevalplus
@@
MBPP+ (Liu et al., 2024c) 378 Python 2023-05 https://huggingface.co/datasets/evalplus/mbppplus
@@
ODEX (Wang et al., 2022e) 945 Python 2022-12 https://huggingface.co/datasets/neulab/odex
@@
ReCode (Wang et al., 2022d) 1,138 Python 2022-12 https://github.com/amazon-science/recode
@@
BigCodeBench (Zhuo et al., 2024a) 1,140 Python 2024-06 https://huggingface.co/datasets/bigcode/bigcodebench
ClassEval (Du et al., 2024) 100 Python 2023-08 https://huggingface.co/datasets/FudanSELab/ClassEval
NaturalCodeBench (Zhang et al., 2024c) 402 Python, Java 2024-05 https://github.com/THUDM/NaturalCodeBench
@@
LiveCodeBench (Naman Jain et al., 2024) 713 Updating Python 2024-03 https://github.com/LiveCodeBench/LiveCodeBench
@@
CRUXEval (Gu et al., 2024) 800 Python 2024-01 https://huggingface.co/datasets/cruxeval-org/cruxeval
@@
Repobench (Liu et al., 2023c) 27k Python, Java 2023-01 https://github.com/Leolty/repobench
@@
  • Repobench (Liu et al., 2023c): Repobench is a benchmark specifically used for evaluating repository-level code auto-completion systems. Supporting both Python and Java, it consists of three interconnected evaluation tasks: RepoBench-R (Retrieval), RepoBench-C (Code Completion), and RepoBench-P (Pipeline).
@@ 5.2. Data Synthesis
Figure 8. The comparison among three representative data synthesis methods used for generating instruction data with LLMs. The Code Alpaca (Chaudhary, 2023) employs the self-instruct method, whereas WizardCoder (Luo et al., 2023) and Magicoder (Wei et al., 2023b) utilize the Evol-Instruct and OSS-Instruct methods, respectively.
Numerous studies have demonstrated that high-quality datasets are integral to enhancing the performance of LLMs in various downstream tasks (Brown, 2020; Meng et al., 2022; Xie et al., 2023; Zhou et al., 2024; Köpf et al., 2024; Wettig et al., 2024). For instance, the LIMA model, a 65B parameter LLaMa language model fine-tuned with a standard supervised loss on a mere 1,000 meticulously curated prompts and responses, achieved performance on par with, or even superior to, GPT-4 in 43% of evaluated cases. This figure rose to 58% when compared to Bard and 65% against DaVinci003, all without the use of reinforcement learning or human preference modeling (Zhou et al., 2024). The QuRating initiative strategically selects pre-training data embodying four key textual qualities — writing style, facts & trivia, required expertise, and educational value — that resonate with human intuition. Training a 1.3B parameter model on such data resulted in reduced perplexity and stronger in-context learning compared to baseline models (Wettig et al., 2024).
Despite these advancements, acquiring quality data remains a significant challenge due to issues such as data scarcity, privacy concerns, and prohibitive costs (Wang et al., 2023a; Liu et al., 2024b). Human-generated data is often labor-intensive and expensive to produce, and it may lack the necessary scope and detail to navigate complex, rare, or ambiguous scenarios. As a resolution to these challenges, synthetic data has emerged as a viable alternative. By generating artificial datasets that replicate the intricacies of real-world information, models such as GPT-3.5-turbo (OpenAI, 2022) and GPT-4 (Achiam et al., 2023) have enabled the creation of rich datasets without the need for human annotation (Wang et al., 2023a; Hämäläinen et al., 2023; Liu et al., 2024b; Laurer, 2024). This approach is particularly beneficial in enhancing the instruction-following capabilities of LLMs, with a focus on generating synthetic instruction-based data.
@@
Drawing on the successes of data synthesis for general-purpose LLMs, researchers have expanded the application of synthetic data to the realm of code generation. The Code Alpaca model, as described in (Chaudhary, 2023), has been fine-tuned on a 7B and 13B LLaMA model using a dataset of 20k instruction-following examples for code generation. This dataset was created by text-davinci-003101010https://platform.openai.com and employed the Self-Instruct technique (Wang et al., 2023a). Building on this, the WizardCoder 15B (Luo et al., 2023) utilizes the Evol-Instruct technique to create an enhanced dataset of 78k evolved code instruction examples. This dataset originates from the initial 20k instruction-following dataset used by Code Alpaca (Chaudhary, 2023), which was also generated by text-davinci-003. The WizardCoder model, fine-tuned on the StarCoder (Li et al., 2023a) base model, achieved a 57.3% pass@1 on the HumanEval benchmarks. This performance not only surpasses all other open-source Code LLMs by a significant margin but also outperforms leading closed LLMs such as Anthropic’s Claude and Google’s Bard. In a similar vein, Magicoder (Wei et al., 2023b) introduces a novel data synthesis approach termed OSS-INSTRUCT which enlightens LLMs with open-source code snippets to generate high-quality instruction data for coding tasks. It aims to address the inherent biases often present in synthetic data produced by LLMs. Building upon CodeLlama (Roziere et al., 2023), the MagicoderS-CL-7B model — fine-tuned with 75k synthetic instruction data using the OSS-INSTRUCT technique and with gpt-3.5-turbo-1106 as the data generator — has outperformed the prominent ChatGPT on the HumanEval Plus benchmark, achieving pass@1 of 66.5% versus 65.9%. In a noteworthy development, Microsoft has introduced the phi-1 model (Gunasekar et al., 2023), a more compact LLM of only 1.3B parameters. Despite its smaller size, phi-1 has been trained on high-quality textbook data sourced from the web (comprising 6 billion tokens) and supplemented with synthetic textbooks and exercises generated with GPT-3.5 (1 billion tokens). It has achieved pass@1 of 50.6% on HumanEval and 55.5% on MBPP, setting a new state-of-the-art for Python coding performance among existing small language models (SLMs). The latest contribution to this field is from the BigCode team, which has presented StarCoder2-15B-instruct (Yuxiang Wei, 2024), the first entirely self-aligned code LLM trained with a transparent and permissive pipeline. This model aligns closely with the OSS-INSTRUCT principles established by Magicoder, generating instructions based on seed functions filtered from the Stack v1 dataset (Kocetkov et al., 2022) and producing responses through self-validation. Unlike Magicoder, StarCoder2-15B-instruct employs its base model, StarCoder2-15B, as the data generator, thus avoiding reliance on large and proprietary LLMs like GPT-3.5-turbo (OpenAI, 2022). Figure 8 illustrates the comparison between Self-Instruct, Evol-Instruct, and OSS-Instruct data synthesis methods.